A light ray strike an air/water surface at an angle of 46 degrees with respect of the normal. find the angle of refraction when the direction of the ray is from the air to water

1
Expert's answer
2022-03-08T08:56:56-0500

The Snell's law says


"n_i\sin\theta_i=n_r\sin\theta _r"

Thus


"\theta_r=\sin^{-1}\left(\frac{n_i}{n_r}\sin\theta_i\right)""\theta_r=\sin^{-1}\left(\frac{1.00}{1.43}\sin46^\circ\right)=30^\circ"